janiya10:

Carter: Carter shared his post with 10 friends, who each share with only 2 people each day. Carter shared his post with 10 friends, who each share with only 2 people each day.

janiya10:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @janiya10 \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle Is there more to the question? \(\color{#0cbb34}{\text{End of Quote}}\) yes ill post it \(\color{#0cbb34}{\text{End of Quote}}\) ok \(\color{#0cbb34}{\text{End of Quote}}\) Amber: Amber shared her photo with 3 people. They continued to share it, so the number of shares increases every day as shown by the function. f(x) = 3(4)x Ben: Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day. Day Number of Shares 0 2 1 6 2 18 Carter: Carter shared his post with 10 friends, who each share with only 2 people each day. Carter shared his post with 10 friends, who each share with only 2 people each da

carmelle:

If im correct, I think the questions are... 1. Write an exponential function to represent the spread of Ben's social media post. 2. Write an exponential function to represent the spread of Carter's social media post. 3. Graph each function using at least three points for each curve. All graphs should be placed together on the same coordinate plane, so be sure to label each curve. You may graph your equation by hand on a piece of paper and scan your work, or you may use graphing technology. 4. Using the functions for each student, predict how many shares each student's post will be received on Day 3 and then on Day 10. Justify your answers. 5. If Amber decides to mail copies of her photo to the 45 residents of her grandmother's assisted living facility, the new function representing her photo shares is f(x) = 3(4)x + 45. How does this graph compare with the original graph of Amber's photo share? 6. Based on your results, which students' post travels the fastest? How is this shown in the equation form of the functions? 7. If you had to choose, would you prefer a post with fewer friends initially but more shares, like Amber, or more friends initially but fewer shares? Justify your answer with your calculations from previous questions

Imagine:

Is it a x-y diagram?

Imagine:

Ah, that's a lot...

janiya10:

Three Algebra 1 students are comparing how fast their social media posts have spread. Their results are shown in the following table. Write an exponential function to represent the spread of Ben's social media post. Write an exponential function to represent the spread of Carter's social media post. Graph each function using at least three points for each curve. All graphs should be placed together on the same coordinate plane, so be sure to label each curve. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. Using the functions for each student, predict how many shares each student's post will be received on Day 3 and then on Day 10. Justify your answers. If Amber decides to mail copies of her photo to the 45 residents of her grandmother's assisted living facility, the new function representing her photo shares is f(x) = 3(4)x + 45. How does this graph compare with the original graph of Amber's photo share? Based on your results, which students' post travels the fastest? How is this shown in the equation form of the functions? If you had to choose, would you prefer a post with fewer friends initially but more shares, like Amber, or more friends initially but fewer shares? Justify your answer with your calculations from previous questions. Amber: Amber shared her photo with 3 people. They continued to share it, so the number of shares increases every day as shown by the function. f(x) = 3(4)x Ben: Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day. Day Number of Shares 0 2 1 6 2 18 Carter: Carter shared his post with 10 friends, who each share with only 2 people each day. Carter shared his post with 10 friends, who each share with only 2 people each day.

Imagine:

My head hurts, I can't..

Imagine:

@supie

carmelle:

Did u do Ben's function already or do u need help with that too?

carmelle:

did u do Bens function?

janiya10:

the module is 4.05

janiya10:

yes i did bens

carmelle:

ok, what did u get?

janiya10:

Ben initially shared his photo with 2 people. And the amount of people who have seen his photo increases by 3 times every day. So for Ben we would have something like: B(x)=2(3)x

carmelle:

ok.. remember that its to the power of x, so it should be 2(3)^x

janiya10:

for ben

carmelle:

yea. for ben

janiya10:

yea

janiya10:

im looking for arter tho what he has

carmelle:

Yea, so Carter started with 10 people so the y- intercept is 10

carmelle:

And it went up by 2 each day, so basically the function would be B(x)=10(2)^x

carmelle:

does that make sense?

janiya10:

yea

janiya10:

thx

carmelle:

yw :)